dodohead06:

jordanJordan has a jar of quarters. When he has 60 quarters in the jar, the jar weighs 52 ounces. When there are 125 quarters in the jar, the jar weighs 65 ounces. What is the weight of the empty jar?

AZ:

60 quarters - the jar weights 52 ounces 125 quarters - the jar weights 65 ounces Obviously the weight is weight of the jar + weight of the quarters We first have to find the weight of the quarters and then we can find the weight of the jar

AZ:

60 quarters = 52 ounces 125 quarters = 65 ounces So do a simple subtraction 125 quarters - 60 quarters = ?? and what would the weight of that many quarters be? 65 ounces - 52 ounces = ?? What do you get?

AZ:

Then we can try to find the weight of 1 quarter and then multiply by 60 to get the weight of 60 quarters and then we subtract that from 52 ounces to find out the weight of an empty jar (since the weight of a jar with 60 quarters is 52 ounces, so if we subtract the weight of 60 quarters then we would get the weight of the empty jar)

crispyrat:

let j be the weight of the jar and q the weight of a quarter 60q+j=52 and 125q+j=65 So using the first equation j=52-60q then just plug it in and then you solve for j